# RewardDS: Privacy-Preserving Fine-Tuning for Large Language Models via Reward Driven Data Synthesis

### v1

@@ Abstract
The success of large language models (LLMs) has attracted many individuals to fine-tune them for domain-specific tasks by uploading their data. However, in sensitive areas like healthcare and finance, privacy concerns often arise. One promising solution is to sample synthetic data with Differential Privacy (DP) guarantees to replace private data. However, these synthetic data contain significant flawed data, which are considered as noise. Existing solutions typically rely on naive filtering by comparing ROUGE-L scores or embedding similarities, which are ineffective in addressing the noise. To address this issue, we propose RewardDS, a novel privacy-preserving framework that fine-tunes a reward proxy model and uses reward signals to guide the synthetic data generation. Our RewardDS introduces two key modules, Reward Guided Filtering and Self-Optimizing Refinement, to both filter and refine the synthetic data, effectively mitigating the noise. Extensive experiments across medical, financial, and code generation domains demonstrate the effectiveness of our method.
@@
Jianwei Wang1, Junyao Yang1, Haoran Li3, Huiping Zhuang1, Cen Chen1,2, Ziqian Zeng1thanks: Corresponding author 1South China University of Technology 2 Pazhou Laboratory, China 3 Hong Kong University of Science and Technology    zqzeng@scut.edu.cn,    wiwjwilliam@mail.scut.edu.cn
@@ 1 Introduction
The remarkable capabilities of Large Language Models (LLMs) in general tasks have motivated many individuals and organizations to customize their own LLMs for domain-specific applications, such as medical diagnosis, financial analysis, etc. (Wu et al., 2023; Chen et al., 2023). While domain adaptation through fine-tuning is attractive, high computational costs make local fine-tuning impractical for most users. Currently, most LLM service providers (Achiam et al., 2023; Yang et al., 2024a; Doubao, 2024) offer fine-tuning services, allowing users to customize LLMs for their needs by preparing and uploading their domain-specific data. However, these data may contain sensitive information, and directly transferring it to the LLM service provider can lead to significant privacy concerns (Zeng et al., 2024; Abdelnabi et al., 2023). Under the client-server context, we consider individuals and organizations seeking to customize LLMs as the clients, the LLM service providers as servers, and the model to be fine-tuned as the target LLM. It remains a critical challenge to develop privacy-preserving fine-tuning methods in such a client-server scenario.
@@
Prior works proposed data synthesis as a promising solution to the challenge (Yue et al., 2023; Kurakin et al., 2023; Yu et al., 2023; Mattern et al., 2022; Flemings and Annavaram, 2024). This approach generates synthetic data to replace the private data used for fine-tuning, thus ensuring privacy protection. Specifically, a generation proxy model is first trained on the private data, optimized by DP-SGD (Abadi et al., 2016) to safeguard privacy. The generation proxy model then generates synthetic data for subsequent LLM training. However, due to the inherent randomness of the sampling process, the synthetic data inevitably contains significant flawed data, including text incoherence or storyline incompleteness, which is considered as noise and leads to less effective LLM fine-tuning, as illustrated in Figure 1 We also provide an example in Figure 7.
@@
Figure 1: Illustration of how RewardDS overcome the dilemma of traditional synthetic data methods. The synthetic data directly sampled from the generation proxy model contain significant flaws, such as incoherent text or incomplete storylines, which are considered noise.
@@
To mitigate the noise, existing methods (Wang et al., 2022; Yu et al., 2024; Xie et al., 2024) proposed to filter out flawed data by measuring its similarity to private data. Wang et al. (2022) use ROUGE-L similarity, while Yu et al. (2024); Xie et al. (2024) compute embedding similarity. However, these metrics fail to evaluate the synthetic data’s effectiveness for domain-specific tasks. Alternative methods (Li et al., 2024b; Wang et al., 2024) sample synthetic data directly from the target LLM on the server to improve quality. But the target LLM is not fine-tuned on domain-specific tasks, so the sampled synthetic data does not help improve model performance on those tasks.
@@
To mitigate the noise in synthetic data while protecting privacy, we propose RewardDS (Reward-driven Data Synthesis), a novel privacy-preserving framework that improves synthetic data quality for the target LLM’s privacy-preserving fine-tuning. RewardDS implements a two-stage quality control process, i.e., filtering and refinement, as illustrated in Figure 1. Specifically, we first train a reward proxy model on private data to assess data quality for domain-specific tasks, using DP-SGD to safeguard privacy. Through Reward Guided Filtering, we apply the reward proxy model to assess synthetic data generated by the generation proxy model and remove samples with low reward scores. Filtering alone may remove a large amount of data, leaving only a small fraction. Therefore, we aim to further refine the synthetic data to obtain more high-quality data. Our Self-Optimizing Refinement module generates multiple candidate responses for each synthetic query and computes their rewards. The generation proxy model analyzes the highest and lowest scoring responses and then generates improvement feedback. Based on this feedback, the target LLM refines the synthetic data following a refinement instruction. The resulting high-quality, filtered, and refined synthetic data is then used to fine-tune the target LLM for domain-specific tasks.
@@ 2 Related Work
2.1 Privacy Preserving Fine-tuning Methods.
Domain-specific data, such as medical diagnoses and financial reports, often contain sensitive information, and directly fine-tuning LLMs on such data raises privacy concerns (Mokhtarabadi et al., 2024; Wang et al., 2023; Jang et al., 2023). Differentially Private Stochastic Gradient Descent (DP-SGD) injects noise into gradients during fine-tuning, ensuring the model does not memorize private data (Abadi et al., 2016; McMahan et al., 2017). Alternatively, data anonymization methods, such as k-anonymity and adversarial anonymization, detect and remove private information to prevent privacy leakage while maintaining model utility (Sweeney, 1997; Romanov et al., 2019; Staab et al., 2024). Another promising approach is generating synthetic data with Differential Privacy (DP) guarantees as a substitute for private data (Yue et al., 2023; Flemings and Annavaram, 2024). This synthetic data is protected by the DP mechanism, contains no user privacy, and can be freely used for further fine-tuning.
2.2 Privacy-Preserving Synthetic Text Generation.
Recent studies have explored synthetic data with differential privacy guarantees as a substitute for private data in LLM fine-tuning, achieving a balance between data utility and privacy protection (Yue et al., 2023; Yu et al., 2024; Kurakin et al., 2023). A series of work has been proposed to reduce computational costs and achieve high-quality synthetic data. Lin et al. (2024); Xie et al. (2024) use APIs and zero-shot learning to generate synthetic data without fine-tuning. Du et al. (2024) combine the strengths of multiple models to generate synthetic data, mitigating the risks associated with relying on a single model. Zou et al. (2025) integrate knowledge from pre-trained language models and generate differentially private synthetic data. Wang et al. (2024) integrates differential privacy with knowledge distillation from professional models, leveraging both local and professional models to generate high-quality synthetic data. However, these methods overlook the noise introduced during the synthetic data sampling process, which can degrade performance. To mitigate this, Yu et al. (2024); Wang et al. (2022) compute the similarity between synthetic and private data to filter out those with low similarity. However, these similarity measures are too surface-level to effectively capture the quality of synthetic data for domain-specific tasks. Therefore, a more robust framework is needed to address the noise in synthetic data and enhance its quality for domain-specific tasks.
@@ 3 Problem Statement
We consider a scenario where the client holds domain-specific data, such as patient’s medical records, which contain sensitive information. Hence, directly transmitting those data to servers for LLM fine-tuning is not allowed. This private data typically is structured as Query-Response pairs, with both query and response containing confidential private information (Wang et al., 2024). The server, which hosts the target LLM, offers only API access while keeping model weights confidential, preventing clients from accessing or locally fine-tuning the model. While clients can fine-tune lightweight LLMs within their computational constraints, these models have inherently weaker capabilities than the target LLMs. This creates a critical challenge: how to leverage a client’s private data to improve the server-hosted LLM’s performance on domain-specific tasks while preserving privacy, given that clients cannot locally fine-tune the target LLM due to inaccessibility of model weights.
Existing methods utilize a lightweight Generation Proxy Model on the client side to generate safe synthetic data for fine-tuning the target LLM on the server (Yue et al., 2023; Yu et al., 2024). However, the randomness of the sampling process may introduce noise in the synthetic data, potentially causing performance degradation. Therefore, our main goal is to explore a more effective method for mitigating the noise in synthetic data, enabling better fine-tuning performance while maintaining user privacy.
@@ 4 Method
To address the performance degradation caused by noise in synthetic data, we propose a novel framework, RewardDS (Reward-driven Data Synthesis), as shown in Figure 2. Our approach additionally trains a Reward Proxy Model on the client side. Then the reward proxy model filters and refines the synthetic data sampled from the generation proxy model through Reward Guided Filtering and Self-Optimizing Refinement modules on the server side. Both modules collaborate to enhance the quality of the synthetic data, driven by the reward signal from the reward model. We will introduce the training process of the generation proxy model and reward proxy model in § 4.1 and the details of reward guided filtering and self-optimizing refinement module are provided in § 4.2.
@@ 4.1 Client Side
The reward model is responsible for evaluating the quality of the synthetic data. It should provide higher rewards for high-quality data while lower rewards for poor-quality data. Following standard reward model training practices Liu et al. (2024), we train the reward proxy model using paired comparison data. Let $W_{0}$ denote the initial backbone model, $W_{gen}$ the fine-tuned generation proxy model, and $W_{\text{rwd}}$ the fine-tuned reward proxy model. For each query $Q$ from the private dataset with its gold response $A_{\text{gold}}$, we generate two responses: $A_{0}$ from $W_{0}$ and $A_{\text{gen}}$ from $W_{gen}$. We then create preference pairs by selecting either $A_{\text{gen}}$ or $A_{\text{gold}}$ as the chosen response $A_{c}$, with $A_{0}$ serving as the rejected response $A_{r}$. The reward proxy model maintains a lightweight architecture for client-side deployment and is fine-tuned using differential privacy (DP-SGD) to prevent privacy leakage.
@@ 4.2 Server Side
We leverage the reward proxy model $W_{\text{rwd}}$ to evaluate each synthetic data and filter out those with low rewards. A lower reward indicates a higher likelihood of the synthetic data being noisy. We select only the top $\lfloor L/k\rfloor$ data, where $L$ is the total number of synthetic data and $k$ is the partition fold (Line 1 in Alg. 1). To compensate for the reduced synthetic dataset size after filtering, we replicate the high-reward data to maintain the total data volume during the target LLM fine-tuning (Line 1 in Alg. 1).
@@
While filtering mitigates noise, it selects only a small subset of samples, potentially leading to overfitting on limited data. Building on LLMs’ self-reflection capabilities (Madaan et al., 2023), we implement a dynamic data refinement strategy to improve low-reward samples, enhancing overall data quality. Initially, for each synthetic query, we generate $N$ candidate responses rather than only one response using the generation proxy model (Line 1 in Alg. 1). The reward proxy model then selects the response with the highest reward score as the chosen response (Line 1 in Alg. 1). We directly fine-tune the target LLM $W_{\text{target}}$ on the chosen response (Line 1 in Alg. 1).
@@
The collaborative process between the reward-guided filtering and self-optimizing refinement modules is presented in Alg. 1. The refinement instruction templates are provided in Appendix H. After the LLM is fine-tuned on the refined synthetic data, it can provide service to the client for those domain-specific tasks.
@@ 5 Privacy Analysis
The only transmitted content between the client and server are the generation proxy model and the reward proxy model. Both models are fine-tuned on the private dataset using the DP-SGD algorithm Abadi et al. (2016). According to the definition of differential privacy (DP) (Dwork and Roth, 2014), adversaries cannot infer any private data from the fine-tuned proxy models. Additionally, based on the post-processing property of the DP framework (Dwork and Roth, 2014), any further operations on the two proxy models will not cause privacy leakage. All subsequent operations on the server, including synthetic data generation, reward-guided filtering, and self-optimizing refinement, are privacy-preserving.
@@ 6.1 Experiments Setup
For the evaluation of the QA task, we employ the ROUGE-1 (R1), ROUGE-L (RL) (Lin, 2004), and Perplexity (PPL) (Hu et al., 2024) as metrics. While automated metrics focus on lexical overlap and fluency, LLM-Judge (Zheng et al., 2023) provides a more comprehensive assessment of semantic accuracy and response quality. Hence, we also use LLM-Judge as a metric. For the code generation task, we use Pass@1 and Pass@10 as evaluation metrics (Chen et al., 2021).
@@
More details on the datasets used and the implementation are provided in Appendix A and Appendix C, respectively.
@@
DP-Generation (Kurakin et al., 2023) fine-tunes the generation proxy model on the client side using DP-SGD. This proxy model is then used to generate synthetic data, which is subsequently utilized to fine-tune the target LLM on the server. DP-Instruct (Yu et al., 2024) introduces additional filtering operations based on the similarity of synthetic queries before LLM fine-tuning; KnowledgeSG (Wang et al., 2024) utilizes the synthetic data to enhance the generation proxy model for domain-specific tasks instead of fine-tuning the target LLM.
More details of the compared method are provided in Appendix B.
@@ 6.3 Main Results
As shown in Table 1, RewardDS outperforms all other baselines across the three domain-specific tasks, except for the PPL on the Medical QA task. DP-Instruct achieves marginally lower PPL in medical QA. This is possibly due to the filtering by similarity, leading the target LLM to overfit on these highly similar samples.
@@
DP-Generation samples synthetic training data to fine-tune the target LLM on the server. However, due to the randomness inherent in the sampling process, the resulting synthetic data contains significant noise, which severely impairs the fine-tuning performance of the LLM on the server. Although DP-Instruct attempts to filter the synthetic data by computing the similarity between the synthetic query and the private query, it still does not perform well. Similarity alone cannot accurately reflect the quality of synthetic data, where higher similarity does not necessarily indicate better data quality.
KnowledgeSG utilizes synthetic data to fine-tune the lightweight proxy model on the server, enhancing it with the assistance of the target LLM. However, the quality of the synthetic data is highly dependent on the target LLM’s capacity for the specific domain task. If the target LLM performs poorly, the synthetic data will likely contain more noise, which could harm the performance of the proxy model. KnowledgeSG performs relatively better on the financial QA task compared to the other two tasks, primarily because the Vanilla LLM performs well on financial QA, whereas it does not on the other tasks. Only in the financial QA task, the performance of the Vanilla LLM surpasses that of the Locally Fine-tuned model, whereas this is not the case for the other tasks.
@@
We also present results after removing the Reward Guided Filtering and Self-Optimizing Refinement modules respectively. We observe the decline in performance across all tasks when either module is removed, which highlights the effectiveness of these modules. Without these modules, more noisy synthetic samples are included during LLM fine-tuning, resulting in performance degradation.
6.4 Evaluation using LLM-Judge
@@
Figure 3: Using LLM-Judge Zheng et al. (2023) to compare the outputs generated by our method with those of other baselines. Win means our method outperformed the baselines, Tie means the results were similar, and Lose means our method performed worse than the baselines.
We also use LLM-Judge (Zheng et al., 2023) for a more reliable evaluation of the medical QA and financial QA tasks. While ROUGE metrics measure lexical similarity to references and PPL captures fluency, these metrics often fail to assess deeper aspects of response quality. Inspired by Zheng et al. (2023), we fine-tune an LLM judger to assess the quality of generated outputs across different baselines. We provide the judger with both the user query and the generated outputs from our method and the baselines, allowing it to determine which is better or declare a tie. Details of the judger training and evaluation process are shown in Appendix D.
As shown in Figure 3, our method outperforms other baselines in both the medical QA and financial QA tasks. DP-Generation and KnowledgeSG struggle with noisy samples from synthetic data, leading to poor performance. Although DP-Instruct filters synthetic data by comparing with private data and removing low-similarity samples, it achieves only limited performance gains compared to DP-Generation. This shows that simple similarity measures do not fully capture the quality of synthetic data. Locally Fine-tuning avoids synthetic data noise by fine-tuning a lightweight proxy model on private data locally, but it still underperforms our method due to the limited learning capacity of the lightweight model for domain-specific knowledge.
6.5 Hyperparameter Analysis
Figure 4: Analysis of hyperparameters including the number of folds $k$ and the number of candidate responses $N$ in Alg. 1 on the medical QA task. To analyze $k$, we set $N=3$; To analyze $N$, we set $k=6$.
We analyze the effect of hyperparameters on our method described in Alg. 1. As shown in Alg. 1, the number of folds $k$ controls the amount of selected synthetic data. A smaller $k$ means more synthetic data is included, but it may also introduce more noise. As illustrated in Figure 4, when $k=1$ (using all the synthetic data), performance decreases. Using a larger $k$ can help exclude noisy data, improving performance. However, setting $k$ too large and excluding too much synthetic data can slightly degrade performance. Therefore, we set $k=6$ for the medical QA task.
The hyperparameter $N$ controls the number of candidate responses. As shown in Figure 4, increasing $N$ slightly improves performance because a larger number of candidates increases the chance of selecting better synthetic data. However, generating more candidates incurs additional time and computational costs. Therefore, for the medical QA task, we set $N=3$.
Further hyperparameter analyses for the legal QA task and code generation task can be found in Appendix E.

### v2

@@ Abstract
The success of large language models (LLMs) has attracted many individuals to fine-tune them for domain-specific tasks by uploading their data. However, in sensitive areas like healthcare and finance, privacy concerns often arise. One promising solution is to generate synthetic data with Differential Privacy (DP) guarantees to replace private data. However, these synthetic data contain significant flawed data, which are considered as noise. Existing solutions typically rely on naive filtering by comparing ROUGE-L scores or embedding similarities, which are ineffective in addressing the noise. To address this issue, we propose RewardDS, a novel privacy-preserving framework that fine-tunes a reward proxy model and uses reward signals to guide the synthetic data generation. Our RewardDS introduces two key modules, Reward Guided Filtering and Self-Optimizing Refinement, to both filter and refine the synthetic data, effectively mitigating the noise. Extensive experiments across medical, financial, and code generation domains demonstrate the effectiveness of our method. Our code and data will be available at https://github.com/wjw136/RewardDS.
@@
Jianwei Wang1, Chengming Shi1, Junyao Yang1, Haoran Li2, Qianli Ma 1, Huiping Zhuang1, Cen Chen1, Ziqian Zeng1thanks: Corresponding author 1 South China University of Technology, 2 HKUST wjwfyu@gmail.com, zqzeng@scut.edu.cn
@@ 1 Introduction
The remarkable capabilities of Large Language Models (LLMs) in general tasks have motivated many individuals and organizations to customize their own LLMs for domain-specific applications, such as medical diagnosis, financial analysis, etc. (Wu et al., 2023; Chen et al., 2023). While domain adaptation through fine-tuning is attractive, high computational costs make local fine-tuning impractical for most users. Currently, most LLM service providers (Achiam et al., 2023; Yang et al., 2024a; Doubao, 2024) offer fine-tuning services, allowing users to customize LLMs for their needs by preparing and uploading their domain-specific data. However, these data may contain sensitive information, and directly transferring it to the LLM service provider can lead to significant privacy concerns (Zeng et al., 2024; Abdelnabi et al., 2023). We denote the individuals or organizations which aim to customize LLMs as client, the LLM service providers as server, and the model being fine-tuned as the target LLM. It remains a critical challenge to develop privacy-preserving fine-tuning methods in such a client-server scenario.
@@
Prior works proposed data synthesis as a promising solution (Kurakin et al., 2023; Yue et al., 2023; Yu et al., 2023; Mattern et al., 2022; Flemings and Annavaram, 2024). These approaches generate synthetic data to replace the private data and can be used for fine-tuning, thus ensuring privacy protection. Specifically, a generation proxy model is first trained on the private data, optimized by DP-SGD (Abadi et al., 2016) to safeguard privacy. The generation proxy model then generates synthetic data for subsequent LLM training. However, due to the inherent randomness of the generation process, the synthetic data inevitably contain significant flawed one, including text incoherence or storyline incompleteness, which is considered as noise and leads to less effective LLM fine-tuning, as illustrated in Figure 1.
@@
Figure 1: Illustration of how RewardDS overcomes the dilemma of traditional synthetic data methods. The synthetic data directly sampled from the generation proxy model contain significant flaws, such as incoherent text or incomplete storylines, which are considered noise.
@@
To mitigate the noise, existing methods (Yu et al., 2024; Wang et al., 2022; Xie et al., 2024) proposed to filter out flawed data by measuring its similarity to private data. Wang et al. (2022) use ROUGE-L similarity, while Yu et al. (2024); Xie et al. (2024) compute embedding similarity. However, these metrics fail to evaluate the synthetic data’s effectiveness for domain-specific tasks. Alternative methods (Wang et al., 2024; Li et al., 2024b) attempt to distill the capabilities from the LLM on the server side into the generation proxy model to support domain-specific tasks. However, since the LLM is not fine-tuned for these specific domains, the distillation provides limited benefit and does not effectively improve task performance. We also present an illustrative example in Figure 7.
@@
To more effectively mitigate noise in synthetic data, we propose RewardDS (Reward-driven Data Synthesis), a novel privacy-preserving framework that improves synthetic data quality for the target LLM’s privacy-preserving fine-tuning. RewardDS implements a two-stage quality control process, i.e., filtering and refinement, as illustrated in Figure 1. Specifically, we first train a reward proxy model on private data to assess data quality for domain-specific tasks, using DP-SGD to safeguard privacy. Through Reward Guided Filtering, we apply the reward proxy model to assess synthetic data generated by the generation proxy model and remove samples with low reward scores. Filtering alone may remove a large amount of data, leaving only a small fraction. Therefore, we aim to further refine the synthetic data to obtain more high-quality data. Our Self-Optimizing Refinement module generates multiple candidate responses for each synthetic query and computes their rewards. The generation proxy model analyzes the highest and lowest scoring responses and then generates improvement feedback. Based on this feedback, the target LLM refines the synthetic data following a refinement instruction. The resulting high-quality, filtered, and refined synthetic data are then used to fine-tune the target LLM for domain-specific tasks.
@@ 2 Related Work
In this section, we will introduce the related work on LLM privacy-preserving fine-tuning methods, which are currently divided into three categories: Anonymity-based methods, Encryption-based methods and Synthesis-based methods.
Anonymity-based methods.
Techniques like k-anonymity and adversarial anonymization can identify and anonymize private data. But they will significantly harm data quality for domain-specific LLM fine-tuning on the server side. (Staab et al., 2024; Sweeney, 1997; Romanov et al., 2019)
Encryption-based methods.
Some approaches employ encryption techniques, such as Homomorphic Encryption (HE) or Secure Multi-Party Computation (SMPC), to protect private data. However, encrypting data and maintaining secure communication between server and client incur significant computational and time overhead, making these methods impractical in real-world scenarios. (Frery et al., 2025; Lou et al., 2020; You et al., 2025)
Synthesis-based methods.
Recent studies have explored using synthetic data with differential privacy (DP) guarantees as a substitute for private data in LLM fine-tuning. While this offers a practical and efficient solution, the synthetic data often contain noisy or flawed samples that significantly hinder LLM fine-tuning. Simple filtering based on text similarity is insufficient to effectively eliminate such noise. (Kurakin et al., 2023; Yue et al., 2023; Yu et al., 2024; Hou et al., 2024; Wang et al., 2024, 2022)
Due to the limited space, a detailed introduction of the above works can be found in Appendix A.
@@ 3 Problem Statement
We consider a scenario where the client holds domain-specific data, such as patient’s medical records, which contain sensitive information. Hence, directly transmitting those data to servers for LLM fine-tuning is not allowed. This private data typically is structured as Query-Response pairs, with both query and response containing confidential private information (Wang et al., 2024). The server, which hosts the target LLM, offers only API access while keeping model weights confidential, preventing clients from accessing or locally fine-tuning the model. While clients can fine-tune some lightweight LLMs within their computational constraints, these models have inherently weaker capabilities than the target LLM. This creates a critical challenge: how to leverage a client’s private data to improve the server-hosted LLM’s performance on domain-specific tasks while preserving privacy.
Existing synthesis-based methods utilize a lightweight Generation Proxy Model to generate safe synthetic data for fine-tuning the target LLM on the server (Yue et al., 2023; Yu et al., 2024). However, the randomness of the generation process introduces significant noise in the synthetic data, potentially causing performance degradation. Therefore, our main goal is to explore a more effective method for mitigating the noise in synthetic data, enabling better fine-tuning performance while maintaining user privacy.
@@ 4 Method
To address the performance degradation caused by noise in synthetic data, we propose a novel framework, RewardDS (Reward-driven Data Synthesis), as shown in Figure 2. Our approach additionally trains a Reward Proxy Model on the client side. Then the reward proxy model filters and refines the synthetic data sampled from the generation proxy model through Reward Guided Filtering and Self-Optimizing Refinement modules on the server side. Both modules collaborate to enhance the quality of the synthetic data, driven by the reward signal from the reward model. We will introduce the training process of the generation proxy model and reward proxy model in §˜4.1. The details of reward guided filtering and self-optimizing refinement module are provided in §˜4.2.
@@ 4.1 Client Side
The reward model is responsible for evaluating the quality of the synthetic data. It should provide higher rewards for high-quality data while lower rewards for poor-quality data. Following standard reward model training practices (Liu et al., 2024), we train the reward proxy model using paired comparison data. Let $W_{0}$ denote the initial backbone model, $W_{gen}$ the fine-tuned generation proxy model, and $W_{\text{rwd}}$ the fine-tuned reward proxy model. For each query $Q$ from the private dataset with its gold response $A_{\text{gold}}$, we generate two responses: $A_{0}$ from $W_{0}$ and $A_{\text{gen}}$ from $W_{gen}$. We then create preference pairs by selecting either $A_{\text{gen}}$ or $A_{\text{gold}}$ as the chosen response $A_{c}$, with $A_{0}$ serving as the rejected response $A_{r}$. The reward proxy model maintains a lightweight architecture for client-side deployment and is fine-tuned using differential privacy (DP-SGD) to prevent privacy leakage.
@@ 4.2 Server Side
We leverage the reward proxy model $W_{\text{rwd}}$ to evaluate each synthetic data and filter out those with low rewards. A lower reward indicates a higher likelihood of the synthetic data being noisy. We select only the top $\lfloor L/k\rfloor$ data, where $L$ is the total number of synthetic data and $k$ is the partition fold (Line 1 in Alg.˜1). To compensate for the reduced synthetic dataset size after filtering, we replicate the high-reward data to maintain the total data volume during the target LLM fine-tuning (Line 1 in Alg.˜1).
@@
While filtering mitigates noise, it selects only a small subset of samples, potentially leading to overfitting on limited data. Building on LLMs’ self-reflection capabilities (Madaan et al., 2023), we implement a dynamic data refinement strategy to improve low-reward samples, enhancing overall data quality. Initially, for each synthetic query, we generate $N$ candidate responses rather than only one response using the generation proxy model (Line 1 in Alg.˜1). The reward proxy model then selects the response with the highest reward score as the chosen response (Line 1 in Alg.˜1). We directly fine-tune the target LLM $W_{\text{target}}$ on the chosen response (Line 1 in Alg.˜1).
@@
The collaborative process between the reward-guided filtering and self-optimizing refinement modules is presented in Alg.˜1. The refinement instruction templates are provided in Appendix J. After the LLM is fine-tuned on the refined synthetic data, it can provide service to the client for those domain-specific tasks.
@@ 5 Privacy Analysis
The only transmitted contents between the client and server are the generation proxy model and the reward proxy model. Both models are fine-tuned on the private dataset using the DP-SGD algorithm Abadi et al. (2016). According to the definition of differential privacy (DP) (Dwork and Roth, 2014), adversaries cannot infer any private data from the fine-tuned proxy models. Additionally, based on the post-processing property of the DP framework (Dwork and Roth, 2014), any further operations on the two proxy models will not cause privacy leakage. All subsequent operations on the server, including synthetic data generation, reward-guided filtering, and self-optimizing refinement, are privacy-preserving. Moreover, we conduct Data Extraction Attack (Carlini et al., 2021) and Membership Inference Attack (Yeom et al., 2018; Choquette-Choo et al., 2021) on our method to empirically demonstrate its privacy protection capability in Section 6.4.
@@ 6.1 Experiments Setup
For the evaluation of the QA task, we employ the ROUGE-1 (R1), ROUGE-L (RL) (Lin, 2004), and Perplexity (PPL) (Hu et al., 2024) as metrics. While automated metrics focus on lexical overlap and fluency, we also use LLM-Judge (Zheng et al., 2023) to provide a more comprehensive assessment of semantic accuracy and response quality. For the code generation task, we use Pass@1 and Pass@10 as evaluation metrics (Chen et al., 2021).
@@
More details on the datasets used and the implementation are provided in Appendix B and Appendix D, respectively.
@@
DP-Generation (Kurakin et al., 2023) fine-tunes the generation proxy model on the client side using DP-SGD. This proxy model is then used to generate synthetic data, which are subsequently utilized to fine-tune target LLM on the server. DP-Instruct (Yu et al., 2024) additionally filters synthetic data based on text similarity before LLM fine-tuning; KnowledgeSG (Wang et al., 2024) distills the capacity from LLM into the generation proxy model to enhance its performance.
More details of the compared method are provided in Appendix C.
@@ 6.3 Main Results
As shown in Table 1, RewardDS outperforms all other baselines across the three domain-specific tasks, except for the PPL on the Medical QA task. DP-Instruct achieves marginally lower PPL in medical QA. This may be attributed to the filtering strategy based on similarity, which could lead the target LLM to overfit on these highly similar samples.
@@
DP-Generation samples synthetic training data to fine-tune the target LLM on the server. However, due to the randomness inherent in the sampling process, the resulting synthetic data contains significant noise, which severely impairs the fine-tuning performance of the LLM on the server. DP-Instruct attempts to filter the synthetic data by computing the similarity between the synthetic query and the private query. But, similarity alone cannot accurately reflect the quality of synthetic data, where higher similarity does not necessarily indicate better data quality. KnowledgeSG distills the capabilities of the target LLM on the server into the generation proxy model for domain-specific tasks. However, since the target LLM is not specifically fine-tuned for these tasks, the improvement through distillation is limited.
@@
We observe consistent performance declines across all tasks when either the Reward Guided Filtering or Self-Optimizing Refinement module is removed, highlighting the importance of both components. Without these components, more noisy synthetic samples are used during LLM fine-tuning, leading to degraded performance.
In addition, following Zheng et al. (2023), we employ an LLM-Judger to assess the generated response from our method and those baseline approaches for QA tasks. Specifically, we provide the LLM-Judger with responses from our method and those from baseline methods, prompting it to judge which response is better. As shown in Figure 3, our method consistently outperforms all baselines on both medical and financial QA tasks. More implementation details are provided in Appendix E. We also provide a case study in Figure 7 and Appendix F to further demonstrate the effectiveness of our method.
@@
Figure 3: Using LLM-Judge to compare the outputs generated by our method with those of other baselines. Win means our method outperformed the baselines, Tie means the results were similar, and Lose means our method performed worse than the baselines.
6.4 Empirical Privacy Protection Results
In this section, we implement Data Extraction Attack (Carlini et al., 2021) and Membership Inference Attack (Yeom et al., 2018; Choquette-Choo et al., 2021) on our method and baselines to empirically evaluate the privacy protection capability. For those baselines, including DP-Generation/DP-Instruct/KnowledgeSG, the client only transfers the generation proxy model to the server. In contrast, RewardDS transfers both the generation proxy model and the reward proxy model. Accordingly, the attack targets for the baselines are limited to the generation proxy model, while for RewardDS, both models can be attacked. We also provide the attack results of No Protection, which serves as the upper bound.
As shown in Figure 4, RewardDS demonstrates superior privacy protection capacity compared to those baselines, as indicated by its comparable ROUGE-L scores and significantly lower F1 scores. This is possibly due to RewardDS allocating the privacy budget for both generation proxy model and reward proxy model, thereby reducing the privacy budget of each individual model and making them more difficult to be attacked.
Implementation details of these attack methods can be found in Appendix G.
Figure 4: Results of Data Extraction Attack and Membership Inference Attack for RewardDS and all baselines under different privacy budgets $\epsilon$ on Medical QA task.